OpenStudy (anonymous):

Which unit rate is the lowest price per ounce? Choice A: 5 ounces of raisins for $1.49 Choice B: 12 ounces of raisins for $3.69

OpenStudy (anonymous):

A. Choice A B. Choice B C. The unit rates are equal. D. The unit rates cannot be determined.

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

It is not c or d.

OpenStudy (anonymous):

I think it is a.

OpenStudy (anonymous):

\[\large \sf \frac{1.49}{5}\] and \[\large \sf \frac{3.69}{12}\] will get you the rates.

OpenStudy (anonymous):

1.49/5 = 0.298

OpenStudy (anonymous):

3.69/12 = 0.3075

OpenStudy (anonymous):

And the smaller one is your answer.

OpenStudy (anonymous):

so, a.

OpenStudy (anonymous):

Yes, it would be a :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem :)